OpenStudy (angel_kitty12):

Is the probability of winning a carnival game was 2/5 and Max played it five times, write an expression that would calculate the probability he won the first three games and lost the last two. Use exponents to express your final answer, but do not evaluate. I don't understand someone please help. Greatly appreciated.

OpenStudy (amistre64):

how would you determine the probability of say: P(AnB) ??

OpenStudy (amistre64):

P(w n w n w n l n l) is what you are determining

OpenStudy (ybarrap):

Let p be probability of winning and q=1-p the probability of losing. You then want the probability of winning 3 followed by 2 losses: \( p\cdot p\cdot p\cdot q\cdot q\). Where p = 2/5 and q = 3/5: \(p^3q^2=\left(\cfrac{2}{5}\right)^3\left(\cfrac{3}{5}\right )^2\)